OpenStudy (anonymous):

The area of a rectangle is 102 square inches and the perimeter is the rectangle is 46 inches. Find the length and width

OpenStudy (anonymous):

Area=Length*Width Perimeter=2*(Width+Legth)

OpenStudy (anonymous):

23=width+length length=width-23 102=(width-23)*width

OpenStudy (anonymous):

then just calculate

OpenStudy (anonymous):

How's you get 23

OpenStudy (campbell_st):

here is the reasoning A = length x width or A = l x w Perimeter = 2length + 2 width so you know 102 = lw 46 = 2l + 2w simplifying the perimeter by dividing by 2 23 = l + w rewrite the equation with l as the subject 23 - w = l now substitute into the area equation 102 = (23 - w)w you have a quadratic to solve.